# url]https://www.bangalab.org

### v1

@@ 3.3.2 Bi-level optimization
The second approach involves reformulating the inner problem using conditions of optimality. This results in a one-level optimization problem (thus, from now on one-level approach). For example, as suggested in [245] in the simpler case of equality constraints on the inner level, first order necessary conditions can be used. A comparison of these approaches on selected benchmark problems can be found in the PhD thesis of Hatz [247]. Simplifying and summarizing, one can expect the two-level approach to work robustly and, to a certain degree, with existing software. However, for larger problems the slow convergence rate will cause a problem. The one-level approach may offer advantages regarding computation time but is more involved from theoretical and practical point standpoints. A reformulation using first order necessary conditions of optimality does not guarantee to find a global optimum and is hindered by the non-smoothness of the complementarity conditions. An overview of approaches and computational aspects of bi-level optimization, involving integrality and uncertainty, can be found in the surveys of [249, 250]. Particularly noteworthy are relaxations and reformulations as mathematical programs with equilibrium constraints [251] and concepts such as lifting [252]. Another alternative for the special case of polynomial systems involves relaxing the inner control problem using a measure problem, which transforms it into a moment problem and transferring it to linear matrix inequalities, as discussed in [253].
@@ 3.3.3 Solving the inner optimization problem
Local solutions to optimal control problems can be calculated using indirect (first-optimize then-discretize) methods. Here the basic idea is to apply the necessary conditions of optimality, Pontryagin’s maximum principle [256], to the optimal control problem. The resulting boundary value problem is then solved with Newton-type methods. One advantage of this approach is that explicit formulae for feedback laws $u^{*}(x(t))$ and formal proofs of optimality might be derived analytically. Alternatively, direct (first-discretize then-optimize) methods rely a finite-dimensional approximation of control functions and path constraints, resulting in a nonlinear programming problem (NLP) in finite dimension. Optimization algorithms are usually iterative. Starting from an initial guess $y^{0}$, they result in iterates $y^{k}$ converging towards a solution of a system of equality and inequality constraints specifying conditions of optimality. A well-known example are the necessary first order Karush-Kuhn-Tucker conditions. As an important advantage, direct methods have been shown to converge for more initial guesses $y^{0}$ when compared to first-optimize then-discretize methods applied to boundary value problems resulting from Pontryagin’s maximum principle. Moreover, they can leverage the sophisticated implementation of established NLP solvers such as ipopt [257] and can be flexibly applied without requiring initial guesses concerning the optimal switching structure. Therefore, we consider direct methods as surveyed in [258, 259, 221], especially direct collocation and direct multiple shooting, as the most promising approaches for general optimal control problems arising on the inner level of (8).
@@
3.3.4 Inference of active constraints
@@ 3.3.6 Summary
The gIOC problem class has high practical relevance with outstanding chances of enabling novel and impactful insights in mathematical biology. At the same time, it has not yet been receiving much attention from the mathematical community at large. We expect that the interesting structure, featuring a data fit problem on the outer and a nonconvex control problem with many unknowns on the inner level, will stimulate cross-discipline research. Example topics are structural properties of the resulting constrained optimization problems, identifiability and observability issues, strategies for designing experiments that yield high-quality data, the development, implementation, and analysis of efficient algorithms, as well as links to symbolic regression and concepts in RL and IRL. We are not aware of any method or solver able to solve it in full generality. One reason for this is that many relevant aspects have only been investigated separately and in different mathematical communities, focussing either on symbolic regression, nonlinear / bi-level / mixed-integer / robust / multi-objective optimization, numerics, optimal control, systems theory, observability and identifiability, mathematical biology, experimental design, or the various research areas associated with machine learning. Even small-scale nonlinear gIOC problems currently remain unsolved, as previous work have focused on closed-loop IOC or on linear dynamics and quadratic costs [264]. However, we are optimistic that with the right interdisciplinary approach, including development of the necessary models, algorithms, theory, and solvers, gIOC can become an enabling technology.

### v2

@@ 3.3.4 Algorithmic approaches to the bi-level problem
The second approach involves reformulating the inner problem using conditions of optimality. This results in a one-level optimization problem (thus, from now on one-level approach). For example, as suggested in [257] in the simpler case of equality constraints on the inner level, first order necessary conditions can be used. A comparison of these approaches on selected benchmark problems can be found in the PhD thesis of Hatz [259]. Simplifying and summarizing, one can expect the two-level approach to work robustly and, to a certain degree, with existing software. However, for larger problems the slow convergence rate will cause a problem. The one-level approach may offer advantages regarding computation time but is more involved from theoretical and practical point standpoints. A reformulation using first order necessary conditions of optimality does not guarantee to find a global optimum and is hindered by the non-smoothness of the complementarity conditions. An overview of approaches and computational aspects of bi-level optimization, involving integrality and uncertainty, can be found in the surveys of [261, 262]. Particularly noteworthy are relaxations and reformulations as mathematical programs with equilibrium constraints [263] and concepts such as lifting [264]. Another alternative for the special case of polynomial systems involves relaxing the inner control problem using a measure problem, which transforms it into a moment problem and transferring it to linear matrix inequalities, as discussed in [265]. As mentioned above, there is also a great number of algorithms that have been suggested in the context of IRL that might be applicable to IOC as well. Particularly interesting might be primal-dual algorithms that iterate between the outer and the inner optimization problems. We refer to the comprehensive IRL surveys [250, 251, 254, 255] for further information.
@@ 3.3.5 Solving the inner optimization problem
Local solutions to optimal control problems can be calculated using indirect (first-optimize then-discretize) methods. Here the basic idea is to apply the necessary conditions of optimality, Pontryagin’s maximum principle [268], to the optimal control problem. The resulting boundary value problem is then solved with Newton-type methods. One advantage of this approach is that explicit formulae for feedback laws $u^{*}(x(t))$ and formal proofs of optimality might be derived analytically. Alternatively, direct (first-discretize then-optimize) methods rely a finite-dimensional approximation of control functions and path constraints, resulting in a nonlinear programming problem (NLP) in finite dimension. Optimization algorithms are usually iterative. Starting from an initial guess $y^{0}$, they result in iterates $y^{k}$ converging towards a solution of a system of equality and inequality constraints specifying conditions of optimality. A well-known example are the necessary first order Karush-Kuhn-Tucker conditions. As an important advantage, direct methods have been shown to converge for more initial guesses $y^{0}$ when compared to first-optimize then-discretize methods applied to boundary value problems resulting from Pontryagin’s maximum principle. Moreover, they can leverage the sophisticated implementation of established NLP solvers such as ipopt [269] and can be flexibly applied without requiring initial guesses concerning the optimal switching structure. Therefore, we consider direct collocation and direct multiple shooting as surveyed in [270, 271, 229] as the most promising approaches for general optimal control problems arising on the inner level of (8).
@@
3.3.6 IOC Extensions
Different extensions are necessary to solve the gIOC problem (8) in full generality.
@@ 3.3.6 IOC Extensions
First steps in the direction of a concurrent inference of objective function and constraints have been taken in the robotics and model predictive control community [276]. Here, often a safe operation is paramount and dealing with a priori unknown obstacles and constraints is practically relevant [277]. An algorithm for learning parametric constraints from locally-optimal demonstrations [278] and the Safe Pontryagin Differentiable Programming paradigm [279] have been suggested. Both leverage necessary conditions of optimality, similar to the one-level approach mentioned in the previous section. A different approach was investigated in [280]. Here the learning of feasible regions is addressed for the special case of linear objectives, using block coordinate descent and mixed-integer linear programming.
The possible switching of objective functions during the observed time horizon has also been addressed in IRL. A multi-phase setting similar to the Brachistochrone example (19) above was considered in [281]. Extending this to a general gIOC setting will involve using theory and algorithms from mixed-integer optimal control [272].
3.3.7 Identifiability and System Excitation
The extension of identifiability concepts to inverse optimal control presents additional layers of complexity beyond traditional system identification. In this context, we aim not only to identify system parameters but also to infer the objective function that explains observed trajectories as optimal solutions to a control problem. For nonlinear ordinary differential equations, this involves establishing conditions under which different objective functions lead to distinguishably different optimal trajectories, i.e. a form of structural identifiability specific to the inverse optimization framework. The challenge is compounded by the fact that optimal trajectories must satisfy necessary conditions from Pontryagin’s Maximum Principle, leading to a two-point boundary value problem where both state and costate variables play crucial roles. Practical identifiability in this context must consider not only measurement noise and sampling limitations but also the possibility that observed trajectories may be only approximately optimal, either due to biological variability or because the true objective function includes terms or constraints not considered in our model class. Furthermore, the nonlinear nature of both the dynamics and the optimization problem can create multiple local optima [162, 57], making it difficult to distinguish between different objective functions that might produce similar trajectories. This necessitates careful consideration of both the model structure and the experimental design to ensure that the observed data contains sufficient information to discriminate between competing hypotheses about the underlying optimization principles governing the biological system.
Persistent excitation is a fundamental concept in system identification that characterizes the richness of input signals required to ensure identifiability [282]. An input signal is persistently exciting when it contains sufficient frequency content to excite all relevant dynamic modes of the system, making them observable in the output data. For linear systems, this is typically expressed through conditions on the spectral content of the input signal or through properties of the information matrix, requiring it to be positive definite.
The role of persistent excitation is to provide sufficient conditions for parameter convergence in identification algorithms, ensuring that the experimental data contains enough information to distinguish between different possible model parameters and structures [283]. Without adequate excitation, multiple parameter sets might explain the observed input-output behavior equally well. While persistent excitation requirements are well-understood for linear systems and can be satisfied by signals such as pseudo-random binary sequences or multi-sine signals, the concept becomes more complex for nonlinear systems, where the required excitation may depend on the operating point and specific nonlinearities involved. This concept bridges the gap between structural and practical identifiability, and is particularly crucial in adaptive control and online identification schemes.
The relationship between persistent excitation, statistical consistency, and identifiability forms a fundamental framework in system identification theory [284]. Persistent excitation ensures that input signals provide sufficient information content, acting as a prerequisite for identifiability. Statistical consistency connects these concepts to the asymptotic behavior of parameter estimates. When an identification method is statistically consistent and the input is persistently exciting, the parameter estimates converge to their true values as the amount of data increases. This requires both structural identifiability (the theoretical possibility of unique parameter identification) and persistent excitation (sufficient information content in the data). The quality of persistent excitation affects the rate of convergence: stronger excitation generally leads to faster convergence and better numerical conditioning of the estimation problem. However, while persistent excitation and structural identifiability are necessary conditions for statistical consistency, they are not always sufficient, particularly in nonlinear systems where issues like local minima and parameter correlation can complicate the relationship between these properties [162]. In practice, these theoretical relationships guide experimental design and help determine how much data is needed to achieve desired estimation accuracy. While this theoretical framework is well-established for engineered systems where inputs can be precisely controlled, its application to biological systems presents unique challenges that we believe warrant careful consideration.
In molecular and cellular biology, achieving persistent excitation is particularly challenging, and in many cases, practically impossible. Biological systems are characterized by complex, nonlinear dynamics with multiple feedback loops, context-dependent responses to stimuli, and inherent stochasticity in cellular processes. These systems typically involve multiple interconnected regulatory networks and homeostatic mechanisms that actively resist perturbations. Furthermore, many system components have limited experimental accessibility, making comprehensive system identification particularly challenging [242].
While modern experimental techniques such as periodic drug administration, optogenetic stimulation, or carefully designed perturbation experiments can help probe system dynamics [285], ensuring complete persistent excitation of all relevant modes remains an open challenge. Moreover, attempts to achieve persistent excitation might push the system outside its normal operating regime, potentially triggering compensatory mechanisms or stress responses that alter the very dynamics we aim to study. It should be noted that the role of noise and stochasticity is crucial, particularly in biological systems where variability arises both from measurement uncertainty and inherent cellular processes.
Overall, identifiability and statistical consistency will play key roles in practical applications. Specifically, multiple experiments increasing data quantity and quality increases the information content and can help overcome these challenges and the conditions under which the inverse problem can be solved despite the practical limitations in achieving persistent excitation. Optimum experimental design (or active learning) is a systematic approach for obtaining useful data. Surveys of the design of optimal experiments in biological processes can be found in [286, 287]. Early works in cell signaling are [288, 289], and for the first time with experimental validation, [290].
Open questions in optimal experimental design (OED) for biological systems revolve around moving beyond traditional parameter refinement to instead handle the inherent uncertainty in these systems. This involves creating methods to distinguish between competing mechanistic models and extract maximum predictive value even from models known to be structurally non-identifiable [288, 291, 292]. The question of whether an observable should be measured at a given time or not results in another layer of mixed-integer optimal control. Compare [293] for a discussion of theoretical properties and numerical methods.
Furthermore, new approaches are needed to define informational richness (the equivalent of persistence of excitation) for biologically feasible inputs, which are often sparse and constrained rather than continuous, and to design experiments that can simultaneously inform dynamics across multiple time scales. The ultimate goal is to develop adaptive, closed-loop OED frameworks where experiments are intelligently updated in real-time to most efficiently resolve these ambiguities, thus accelerating scientific discovery. Designing optimal experiments not only for parameter estimation but also for the concurrent discrimination of objectives, constraints, and models is completely open.
@@ 3.3.9 Summary
The gIOC problem class has high practical relevance with outstanding chances of enabling novel and impactful insights in mathematical biology. At the same time, it has not yet been receiving much attention from the mathematical community at large. We expect that the interesting structure, featuring a data fit problem on the outer and a nonconvex control problem with many unknowns on the inner level, will stimulate cross-discipline research. Example topics are structural properties of the resulting constrained optimization problems, identifiability and observability issues, strategies for designing experiments that yield high-quality data, the development, implementation, and analysis of efficient algorithms, as well as links to symbolic regression and concepts in RL and IRL. We are not aware of any method or solver able to solve it in full generality. One reason for this is that many relevant aspects have only been investigated separately and in different mathematical communities, focussing either on symbolic regression, nonlinear / bi-level / mixed-integer / robust / multi-objective optimization, numerics, optimal control, systems theory, observability and identifiability, mathematical biology, experimental design, or the various research areas associated with machine learning. Currently, the most active community in method development seems to be related to robotics and reinforcement learning as discussed in Sections 3.3.3 and 3.3.4. We are optimistic that with the right interdisciplinary approach, including development of the necessary models, algorithms, theory, and solvers, gIOC can become an enabling technology also in molecular and cell biology.
@@ 4 Conclusion
Acknowledgements. JRB acknowledges financialsupport from grant PID2020-117271RB-C22 (BIODYNAMICS) funded by MCIN/AEI/10.13039/501100011033, from grant PID2023-146275NB-C22 (DYNAMO-bio) funded by MICIU/AEI/ 10.13039/501100011033 and ERDF/EU, and from grant CSIC PIE 202470E108 (LARGO). SS acknowledges financial support from the research initiative “SmartProSys: Intelligent Process Systems for the Sustainable Production of Chemicals” funded by the Ministry for Science, Energy, Climate Protection and the Environment of the State of Saxony-Anhalt and from the German Research Foundation DFG via grant 314838170, GRK 2297 MathCoRe.
We thank Eva Balsa-Canto, Steffen Klamt, Mariko Okada, Andreas Schuppert, and Alejandro F. Villaverde for stimulating discussions and feedback.